OpenStudy (anonymous):

Solve the equation for y. 7x - 3y = 4

OpenStudy (nottim):

nothing else provided?

OpenStudy (nottim):

if so, all you can do iss 'rearrange' for y...

OpenStudy (anonymous):

Use substitution.

OpenStudy (anonymous):

7(0) - 3y = 4

OpenStudy (anonymous):

7x - 3(0) = 4

OpenStudy (anonymous):

Then solve. fractions are not forbidden

OpenStudy (anonymous):

it wants me to figure out what y is but i don't know how to solve it...answer is either y = x- or y = x+

OpenStudy (anonymous):

Okay then. 7(0) - 3y = 4 -3y = 4 y = 4/-3

OpenStudy (anonymous):

7x - 4/-3 = 4 7x = 16/3 Divide both by 7 x = 16/3 *1/7 = 16/21 So x = 16/21 and y = 4/-3